### v1

# Z-Dip: a validated generalization of the Dip Test

Abstract
Detecting multimodality in empirical distributions is a fundamental problem in statistics and data analysis, with applications ranging from clustering to social science. Hartigan’s Dip Test is a classical nonparametric procedure for testing unimodality versus multimodality, but its interpretation is hindered by strong dependence on sample size and the need for lookup tables. We introduce the Z-Dip, a standardized extension of the Dip Test that removes sample-size dependence by comparing observed Dip values to simulated null distributions. We calibrate a universal decision threshold for Z-Dip via simulation and bootstrap resampling, providing a unified criterion for multimodality detection. In the final section, we also propose a downsampling-based approach to further mitigate residual sample-size effects in very large datasets. Lookup tables and software implementations are made available for efficient use in practice.
1 Introduction
The ability to test whether a distribution is unimodal or multimodal is crucial across many domains. In clustering, multimodality indicates the presence of distinct groups. In social and political science, multimodality in opinion distributions is often interpreted as evidence of polarization. Statistical tests of modality thus provide an essential link between empirical data and theoretical interpretation.
Hartigan and Hartigan’s Dip Test [1] remains one of the most widely used nonparametric methods for testing unimodality, and has found applications across diverse fields [2, 3, 4]. The test statistic, known as the Dip statistic, measures the maximum deviation between the empirical distribution function and the closest unimodal distribution function. Critical values and $p$-values under the null hypothesis of unimodality are determined by simulation under a uniform distribution.
Despite its popularity, the Dip Test suffers from three major limitations. First, its critical values are strongly dependent on sample size, requiring large lookup tables or resimulation for each application. Second, the test statistic is not standardized, preventing meaningful comparisons of the strength of multimodality across datasets. Third, for large samples, the Dip Test tends to reject unimodality even for nearly unimodal distributions, leading to an undesirable loss of resolution.
In this paper, we propose a simple yet effective generalization of the Dip Test, which we call the Z-Dip. By standardizing the observed Dip against its null distribution, the Z-Dip provides a scale-free measure of deviation from unimodality that can be compared across sample sizes. Moreover, we calibrate a single universal threshold for statistical significance, eliminating the need for sample-size–specific lookup tables. The resulting algorithm, implemented in R and distributed alongside this paper, retains the same $O(n\log n)$ computational complexity as the original Dip Test while offering more consistent and interpretable results.
We validate the Z-Dip both on simulated samples, and on an empirical sample of more than 117,000 opinion distributions, finding effectively complete agreement between the proposed Z-Dip threshold and the Dip Test’s $p$-value decisions. Finally, we address the issue of information loss in large samples and introduce an alternative correction, which enables consistent comparisons across large datasets.
By explicitly characterizing how sample size affects standardized multimodality statistics, and by offering a transparent correction procedure, this work aims to improve the interpretability and comparability of such metrics. The correction can be implemented with minimal changes to existing workflows and with virtually no added computational overhead. In this sense, the simplicity of the approach is its main strength: it provides a clear and general remedy to a problem that is otherwise difficult to detect.
2 Background
2.1 Hartigan’s Dip Test
Hartigan’s Dip Test, introduced by Hartigan and Hartigan in 1985 [1], is a classical nonparametric method for testing whether a sample distribution is unimodal or multimodal. Formally, it tests the null hypothesis of unimodality against the alternative of multimodality, using the uniform distribution as the “least favorable unimodal null”. The test quantifies how far the empirical cumulative distribution function (ECDF) of the data deviates from its respective closest unimodal cumulative distribution function: intuitively, a unimodal distribution is one whose density has a single peak, which implies that its cumulative distribution is convex up to the mode and concave thereafter. The Dip Test exploits this shape property by finding the unimodal CDF that minimizes the maximum vertical distance from the observed ECDF. This minimum distance, known as the Dip statistic, summarizes how strongly the observed distribution departs from unimodality. In essence, the Dip measures the smallest possible maximum discrepancy between the observed data and any unimodal distribution. Larger Dip values indicate greater deviation from unimodality, and therefore stronger evidence of multimodality.
The sampling distribution of the Dip under the null hypothesis is obtained by simulation from the uniform distribution on $[0,1]$, which represents the boundary case of a unimodal distribution. Hartigan and Hartigan provided tabulated critical values for a range of sample sizes, allowing the test to be implemented via lookup or resimulation. A key strength of the Dip Test lies in its simplicity: it is fully nonparametric and requires no user-specified parameters, bandwidths, or smoothing assumptions. Unlike kernel- or mixture-based methods, the Dip Test can be applied “out of the box,” with no tuning or model selection. This property has contributed to its wide adoption, as researchers can directly apply it to empirical data without additional calibration or subjective choices. Thanks to its aforementioned simplicity and robustness, since its introduction, the Dip Test has been widely adopted across fields [5, 2, 3, 4].
2.2 Limitations of the Classical Dip Test
Despite its merits, the classical Dip Test presents several practical and conceptual limitations that hinder its interpretability and use in large-scale applications.
  1. (i)
    Dependence on sample size. The null distribution of the Dip statistic varies strongly with sample size $N$. As $N$ increases, the Dip values expected under unimodality decrease systematically, since larger samples provide finer-grained approximations to the true distribution. Consequently, each sample size requires its own set of critical values or a separate Monte Carlo calibration. This dependence complicates both interpretation and automation, as results from samples of different sizes cannot be directly compared.
  2. (ii)
    Lack of standardization. The Dip statistic itself is not standardized. Its raw magnitude depends on the scale of sampling variability and on $N$, making it unsuitable for quantitative comparison across datasets. For example, a Dip of $0.1$ may signal multimodality in a small sample, but be negligible in a large one. Without a standardized scale, the test can only yield binary decisions (reject or not reject unimodality), rather than interpretable effect sizes that quantify “how multimodal” a distribution is. However, the Dip statistic already encapsulates this information: it measures the strength of deviation from unimodality, but this signal is obscured by sample-size dependence.
  3. (iii)
    Inconsistencies at large sample sizes. At large $N$, the Dip Test becomes increasingly sensitive to even minor deviations from the ideal unimodal shape. Because the null distribution of the Dip shrinks toward zero as $N$ grows, the smallest irregularities, such as sampling noise or slight skewness, are sufficient to reject unimodality. In practice, this produces false positives in large samples: distributions that are visually unimodal but statistically classified as multimodal. This issue means that the Dip Test’s binary decision rule becomes less informative with increasing $N$, effectively turning every large sample into a “significant” case. As a result, the test loses its ability to distinguish meaningful multimodality (e.g., distinct peaks) from trivial numerical deviations.
Together, these limitations point to the need for a standardized, scale-free reformulation of the Dip Test that preserves its parameter-free simplicity while allowing meaningful comparison across sample sizes, maintaining interpretability at large $N$, and eliminating the need for case-specific calibration. The next section introduces the Z-Dip, which addresses these issues by standardizing the Dip relative to its null distribution.
3 Z-Dip
3.1 Definition of the Z-Dip Statistic
We propose a standardized version of Hartigan’s Dip statistic, which we call the Z-Dip. For an observed dataset of size $N$, let $Dip_{obs}$ denote its Dip statistic’s value. We define
$Z\text{-}Dip=z=\frac{Dip_{obs}-\mu_{N}}{\sigma_{N}},$ (1)
where $\mu_{N}$ and $\sigma_{N}$ denote the mean and standard deviation of Dip values under the null distribution of $N$-sized samples drawn from the uniform distribution on $[0,1]$.
The score proposed in Formula 1 quantifies how strongly the observed distribution deviates from unimodality, in units of standard deviation relative to the null expectation. Positive values indicate stronger evidence of multimodality, while values near zero or negative indicate conformity with unimodality. Because it rescales the Dip relative to its null, the Z-Dip provides a scale-free, comparable measure across different sample sizes.
3.2 Simulation of Null Distributions
To estimate $\mu_{N}$ and $\sigma_{N}$, we simulate large ensembles of null samples from the uniform distribution for a grid of sample sizes $N$ ranging from small ($N=4$) to very large (N=72, [^0] N=72{,}000). The uniform distribution is used in line with Hartigan and Hartigan’s original formulation, where it was used to calibrate empirical percentage points of the dip statistic. For each simulated dataset, we compute the Dip statistic, yielding an empirical null distribution. The mean and standard deviation of this distribution provide estimates of $\mu_{N}$ and $\sigma_{N}$.
To enable efficient application, these values are stored in precomputed lookup tables. For sample sizes not explicitly simulated, we interpolate $\mu_{N}$ and $\sigma_{N}$ linearly between adjacent grid points. This precomputation and interpolation ensure that Z-Dip scores can be rapidly obtained for any sample size without repeated simulation, and the binary file containing the complete table is provided with the software accompanying this paper.
4 Results
4.1 Comparison with the Classical Dip Statistic
We first compare the behavior of the raw Dip and the standardized Z-Dip across sample sizes. The classical Dip values show strong sample-size dependence: as $N$ increases, the expected Dip under the null shrinks toward zero, while its variance decreases. This makes large-sample values numerically small and difficult to interpret.
By contrast, the Z-Dip eliminates this dependence. Across all simulated sample sizes, the null distribution of the Z-Dip is centered around zero with approximately constant spread (Figure 1, panel A). This confirms that standardization removes the sample-size dependence inherent in the raw Dip statistic.
A direct comparison between Dip $p$-values and corresponding Z-Dip values (Figure 1, panel C) further demonstrates the strong monotonic relationship between the two measures, in contrast with the regular Dip statistic (Figure 1, panel B). This alignment suggests that a single Z-Dip threshold can reproduce conventional significance levels across all $N$.
Figure 1: Comparison of the classical Dip statistic and the standardized Z-Dip across sample sizes. Z-Dip distributions were obtained over 9,999 simulated uniform samples, with 1,000 iterations of Z-Dip computation for each simulated sample. (A) Density of Z-Dip scores for different sample sizes (NN), showing strong overlap of null distributions, proving that standardization removes the dependence on $N$. (B, C) Scatterplot of Dip statistic and Z-Dip values versus corresponding Dip Test $p$-values for simulated unimodal distributions, colored by sample size $N$. The strong monotonic relationship indicates that a single Z-Dip threshold consistently matches the $p<0.05$ level across all $N$.
4.2 Threshold Calibration
Under the null hypothesis of unimodality, the Z-Dip is approximately normally distributed with mean zero and unit variance. If this normal approximation were exact, a two-tailed 5% significance level would correspond to the canonical threshold $z=1.96$.
However, empirical simulations reveal that the null distribution of the Z-Dip values is slightly right-skewed (Fig. 1, panel A). Even after standardization, minor deviations from normality persist, likely due in part to the numerical estimation of $\mu_{N}$ and $\sigma_{N}$. As a result, adopting the theoretical $z=1.96$ cutoff could yield a false-positive rate marginally above the nominal 5% level.
To ensure robustness across sample sizes, we estimate the Z-Dip threshold corresponding to $p=0.05$ empirically. For each $N$, we compute Dip-based $p$-values and corresponding Z-Dip scores for a large ensemble of null samples, then identify the Z-Dip value at which $p=0.05$. Bootstrap resampling provides confidence intervals for each estimated threshold. These values are highly consistent across $N$, with upper bounds ranging narrowly around 1.97. We therefore adopt a conservative universal cutoff $z=1.975$, corresponding closely to the theoretical $1.96$ threshold while ensuring that the nominal Type I error rate is not exceeded for any sample size.
4.3 Computational Considerations
The computation of the Dip statistic, as originally defined by Hartigan and Hartigan, involves finding the maximum deviation between the empirical distribution function and the closest unimodal distribution. The algorithm proceeds by sorting the observed data, constructing candidate unimodal envelopes, and computing the maximal deviation. The sorting step dominates the computational cost, leading to an overall time complexity of $O(n\log n)$. After sorting, the iterative identification of modal intervals and the computation of cumulative deviations are linear in $n$, so the asymptotic complexity is determined by the sorting operation.
The Z-Dip retains this computational structure. Once the Dip statistic $Dip_{obs}$ is computed, the standardized Z-dip score, defined in Formula 1, is obtained by a constant number of arithmetic operations, which have $O(1)$ complexity. Therefore, the total runtime of the Z-Dip remains $O(n\log n)$.
Crucially, the Z-Dip does not require repeated Monte Carlo sampling or permutation testing to obtain significance levels. Once lookup tables for $\mu_{N}$ and $\sigma_{N}$ are precomputed, evaluating the Z-Dip for new samples becomes effectively instantaneous. The absence of computational overhead makes the Z-Dip suitable for high-throughput or large-scale applications, such as distributional analysis across thousands of subpopulations or time points.
4.4 Synthetic Testing
We validated the Z-Dip statistic on controlled synthetic distributions with known modality structure. Synthetic samples were generated using a Gaussian mixture simulator of the form
$y_{i}\sim\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}(\mu_{k},\sigma_{k}^{2}),$
where $K$ is the number of modes, and
$\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{K}),\quad\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{K}),\quad\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{K})$
denote respectively the vectors of means, standard deviations, and mixing proportions of the $K$ Gaussian components, with $\pi_{k}\geq 0$ and $\sum_{k=1}^{K}\pi_{k}=1$.
The generator function allows full control over the number, separation, and spread of modes, enabling systematic testing from clearly unimodal to strongly bimodal or multimodal cases. For each configuration, we simulated multiple sample sizes and computed both the classical Dip Test $p$-values and the corresponding Z-Dip scores.
Across all simulations, the Z-Dip preserved the inferential behavior of the Dip Test, with perfect correspondence between significance decisions.
Table 1 summarizes the results across unimodal, weakly bimodal, strongly bimodal, and trimodal conditions. The specific data-generating processes were defined as follows:
  • Unimodal: ${\mu}=(-0.3,-0.3)$, ${\sigma}=(0.1,0.1)$, ${\pi}=(1,0)$
  • Weakly bimodal: ${\mu}=(-0.6,0.6)$, ${\sigma}=(0.15,0.15)$, ${\pi}=(0.7,0.3)$
  • Strongly bimodal: ${\mu}=(-0.6,0.6)$, ${\sigma}=(0.1,0.1)$, ${\pi}=(0.5,0.5)$
  • Trimodal: ${\mu}=(-0.75,0,0.75)$, ${\sigma}=(0.1,0.1,0.1)$, ${\pi}=(1/3,1/3,1/3)$
Multimodal mixtures yielded large positive Z-Dip values and were consistently detected as such, while unimodal distributions remained centered near zero or negative values. It can also be noted that, for genuinely multimodal distributions, the Z-Dip tends to grow systematically with sample size $N$. This behavior arises because well-separated modes produce a persistent “dip” in the empirical cumulative distribution function: as $N$ increases, the separation between peaks is resolved more clearly, leading to larger Dip statistics and correspondingly higher Z-Dip scores. This behavior is further illustrated in Fig. S1 of the Supplementary Information. In contrast, unimodal distributions exhibit flat or slowly varying Z-Dip values with $N$, since noise dominates and no persistent secondary mode exists.
Table 1: Synthetic validation of the Z-Dip statistic under varying modality and sample size (1000 replications per condition).
Configuration NN Mean Z-Dip $p<0.05$ (Dip Test) $Z>1.975$ Agreement
Unimodal 50 -0.86 0.01 0 0.998
100 -1.04 0.00 0 1.000
200 -1.22 0.00 0 1.000
400 -1.41 0.00 0 1.000
500 -1.47 0.00 0 1.000
Weak bimodal 50 7.16 1.00 1 1.000
100 11.38 1.00 1 1.000
200 17.20 1.00 1 1.000
300 21.38 1.00 1 1.000
400 25.93 1.00 1 1.000
500 28.93 1.00 1 1.000
Strong bimodal 50 9.64 1.00 1 1.000
100 15.16 1.00 1 1.000
200 22.55 1.00 1 1.000
300 28.06 1.00 1 1.000
400 33.78 1.00 1 1.000
500 37.86 1.00 1 1.000
Trimodal 50 4.48 1.00 1 0.999
100 7.10 1.00 1 1.000
200 10.70 1.00 1 1.000
300 13.34 1.00 1 1.000
400 16.19 1.00 1 1.000
500 18.12 1.00 1 1.000
These synthetic results provide a controlled validation, showing that the Z-Dip behaves as theoretically expected under both null (unimodal) and alternative (multimodal) conditions.
4.5 Empirical Validation
To evaluate the empirical validity of the proposed Z-Dip statistic, we applied it to a large-scale dataset of 117,457117{,}457 observed opinion distributions. Each distribution represents the political leaning of YouTube users engaging in political video discussions. User leanings are continuous in the range [1,+1][-1,+1], where 1-1 indicates strongly left-leaning and +1+1 strongly right-leaning users. These distributions were previously constructed and described in [4], and provide a rich and heterogeneous collection of empirical distributions, enabling large-scale testing of modality detection methods. For each distribution, we computed both the classical Dip statistic and the standardized Z-Dip, using the universal threshold $z=1.975$ for significance. Despite the empirical sample being heavily skewed toward unimodal (i.e., non-multimodal) cases, the two methods produced nearly identical classifications, with an overall agreement of $99.63\%$, as shown in Fig. 2. This close correspondence confirms that the Z-Dip effectively reproduces the inferential behavior of the original Dip Test while providing a standardized, comparable measure of deviation from unimodality.
The few discrepancies (450 cases in total) occur almost exclusively at the lowest sample sizes. The median number of observations among the disagreeing cases is $N=11$, while the third quartile is $N=52$. This pattern seems consistent with small-sample limitations of the original Dip Test: when $N$ is small, the empirical distribution contains too few distinct values to approximate a smooth shape, leading to discrete artifacts and unstable estimates. Under these conditions, even minimal numerical variation, such as rounding differences or the interpolation of $\mu_{N}$ and $\sigma_{N}$ in the Z-Dip, can flip the classification outcome.
It is also worth noting that a fixed universal threshold ($z=1.975$) necessarily ignores the slight $N$-dependence of the Dip’s true critical values. Minor disagreements are therefore expected in boundary cases, particularly for very small samples. Nonetheless, across realistic sample sizes, the agreement between the Z-Dip and the classical test is effectively perfect.
In sum, these results demonstrate that the Z-Dip not only preserves the decision structure of the Hartigan Dip Test but also generalizes it into a form that is computationally efficient, interpretable, and robust across datasets of varying sizes. In practice, the Z-Dip can be treated as a drop-in replacement for the classical Dip Test that “just works”, eliminating the need for sample-size-specific calibration without sacrificing accuracy or interpretability.
Figure 2: Confusion matrix of the agreement between the Z-Dip threshold and the Dip Test’s $p$-value over the 117,457117{,}457 empirical samples considered.
5 Addressing Large-Sample Sensitivity
A limitation of the classical Dip Test, and by extension the Z-Dip, is its increasing sensitivity at very large sample sizes. As $N$ grows, the null distribution of the Dip statistic contracts toward zero, so even minor irregularities (such as slight skewness or minimal outliers) can produce statistically significant deviations. This leads to a proliferation of “false positives”, where visually unimodal distributions are classified as multimodal. While the Z-Dip standardization improves comparability across sample sizes, the fundamental sensitivity to tiny deviations in large-$N$ samples remains.
One straightforward solution is to resample or downsample the data to a moderate effective sample size before computing the Dip statistic or the corresponding Z-Dip. By reducing $N$ to a size where the null distribution retains meaningful variance, the test becomes less sensitive to trivial deviations while preserving the ability to detect substantive multimodality. Concretely, for a large sample of size $N$, one can compute the Z-Dip for $n_{sim}$ times, with $n_{sim}$ large, on random subsets of size $N_{\text{sub}}\ll N$, and then average the resulting Z-Dip scores:
$\overline{Z\text{-Dip}}=\frac{1}{n_{\text{sim}}}\sum_{i=1}^{n_{\text{sim}}}Z\text{-Dip}(x_{i}),$
where $x_{i}$ is a random subset of size $N_{\text{sub}}$, and $Z\text{-Dip}(x_{i})$ is the Z-Dip test computed on the subset $x_{i}$. This procedure preserves the nonparametric, tuning-free nature of the test, while mitigating the pathological sensitivity of the statistic in very large samples.
While marginally more computationally intensive, this downsampling-based approach enables meaningful comparisons across very large datasets and ensures that the test captures substantive deviations from unimodality rather than trivial sampling noise.
Figure 3 illustrates this effect: for two visually similar distributions, one with N=1, [^0] N=1{,}000 and another with N=100, [^0] N=100{,}000, with both having a negligible secondary peak further from the mode, the classical Dip Test, along with its $p$-value and the corresponding Z-Dip score, classify the smaller sample as unimodal, and the larger sample as multimodal. In contrast, the downsampled Z-Dip correctly identifies both distributions as unimodal ($z<1.975)$, demonstrating that it effectively controls for scale-dependent artifacts and preserves meaningful inference about the true modality of the data.
Figure 3: Effect of sample size on Dip Test results and the corresponding downsampled Z-Dip correction. Both samples in the figure contain a detectable but practically negligible secondary mode around $X=0.5$, which leads the test to classify the large sample as multimodal. The downsampled Z-Dip, however, correctly identifies the distribution as unimodal, illustrating its robustness to scale-induced artifacts.
While simple, this procedure remains stable across sample sizes and distributional conditions. To verify this, we conducted a simulation using 132 approximately log-spaced sample sizes, ranging from $N=150$ to N=72, [^0] N=72{,}000, to ensure uniform coverage on a multiplicative scale. Samples were drawn from the same underlying distributions (unimodal, bimodal, and uniform). For each case, we repeatedly computed the downsampled Z-Dip using fixed subsamples of size $N_{\text{sub}}=100$. Even for a small number of 30 iterations of subsampling, which adds little computational overhead, the resulting values remained highly consistent across all tested $N$, indicating that the downsampling-based correction neutralizes the scale-dependent inflation observed in the classical Dip and Z-Dip statistics. Figure 4 illustrates this stability, showing that the downsampled Z-Dip values fluctuate minimally even as the total sample size increases by orders of magnitude.
Figure 4: Stable trend of downsampled Z-Dip for sample sizes $N$ ranging from 150 to 72,000, computed for data drawn from three distributions: (i) a bimodal Gaussian mixture with $\mu=(-0.6,0.6)$, $\sigma=(0.1,0.1)$, and equal component weights; (ii) a weakly bimodal distribution with $\mu=(-0.6,0.6)$, $\sigma=(0.1,0.2)$, and a minor mode weight of $0.025$; and (iii) a uniform distribution on $[0,1]$. Each point represents the mean downsampled Z-Dip over $n_{sim}=30$.
In principle, other strategies could also be applied, such as smoothing the ECDF before computing the Dip, or incorporating a sample-size-adjusted penalty term. We leave a systematic exploration of these extensions for future work, but note that even the simple downsampling approach is sufficient to restore interpretability in ambiguous, high-$N$ scenarios.
6 Conclusions
This work introduced the Z-Dip, a standardized reformulation of Hartigan’s Dip Test that enables consistent and interpretable multimodality assessment across sample sizes. By expressing the Dip in standardized units relative to its null distribution, the Z-Dip transforms a size-dependent hypothesis test into a scale-free descriptive measure. The proposed framework retains the theoretical foundation and computational efficiency of the original Dip Test, while eliminating the need for lookup tables and enabling direct comparison of multimodality strength across datasets.
Validation on synthetic samples and empirical data demonstrate that the Z-Dip preserves the inferential behavior of the Dip Test, achieving virtually perfect correspondence in significance decisions. At the same time, the standardized scale provides additional interpretability, allowing researchers to quantify how strongly a distribution deviates from unimodality.
Finally, we proposed a downsampling-based correction to stabilize Z-Dip values in very large samples, enabling robust comparisons across datasets of varying size. Together, these contributions establish the Z-Dip as a practical, transparent, and computationally efficient tool for detecting and quantifying multimodality. We believe that the availability of reference tables and open-source implementations will facilitate its adoption in applications in a wide range of fields.
Acknowledgements
We thank Parnasso Nesso for their precious inspiration and contribution.
Supplementary Information
Power-law like scaling of Z-Dip for multimodal distributions
As introduced in Table 1 of the main manuscript, for multimodal samples generated from the same underlying distribution, Z-Dip values follow an approximately power-law behavior as a function of $N$ (Figure S1), such that:
$E[\text{Z-Dip}(N)]\propto N^{\alpha}$
Figure S1: Scaling of Z-Dip as NN grows. The average Z-Dip values were computed over 100 samples coming from the same underlying distribution, for each NN.